OpenStudy (rob1525):

The height of a box is one more than its length l, and the length is one more that its width w. Find a polynomial for the volume V in terms of the following: the width w, the length l, and the height h.

OpenStudy (ddcamp):

What is the formula for volume?

OpenStudy (rob1525):

V= l*w*h

OpenStudy (rob1525):

Ive got h= L + 1, L= w + 1, W = L -1

OpenStudy (ddcamp):

OK. If h = l+1, and w=l-1, what would L*w*h equal?

OpenStudy (ddcamp):

L*(L+1)*(L-1) = ?

OpenStudy (rob1525):

oh wow, so (L^2 +L) (L-1) = L^3-L^2 + L^2 -L = L^3-L

OpenStudy (rob1525):

so this is in terms of L right?